It takes you 4 days to paint a house if you paint for eight hours a day. It takes your brother 5 days. How long would it take you to pain the house together? Answers should be in days, hours, minutes form.
Please show the work, I have the answer but don't know how to figure it out.

Answer by scott8148(6628)

working in hours is probably easiest

each person paints a FRACTION of the house based on their individual time and the combined time (t)
___ the fractions sum to one (the WHOLE house)

[t / (4*8)] + [t / (5*8)] = 1

5t + 4t = 160 ___ t = 160/9 hours